OpenStudy (anonymous):

write an equation for the inverse f(x)= x-7 ----- 2

OpenStudy (anonymous):

what do you mean by mutliply both sides by 2?

OpenStudy (turingtest):

1) rename f(x) and call it y 2) multiply both sides by 2 3) add 7 to both sides 4) switch the places of x and y 5) rename y f^(-1)(x) multiply both sides by 2 is like so: x=y/2 2x=y

OpenStudy (anonymous):

so the answer wouldnt be y=2x-14/4 after i multpily both sides by 2?

OpenStudy (turingtest):

no, the 2's on the right cancel you also forgot to multiply the right side:\[y=\frac{x-7}2\]\[2y=x-7\]now continue with the rest of the steps

OpenStudy (turingtest):

I meant you forgot to multiply the left side*

OpenStudy (anonymous):

okay so now i add 7 to -7 and x right? or to the 2y? im so confused on this stuff!

OpenStudy (turingtest):

you add it to /both/ sides in an equation, if you do something to one side you must always do it to the other as well, or else the two sides won't be equal anymore

OpenStudy (anonymous):

alright, that makes sense.

OpenStudy (turingtest):

good :D so go ahead and give it a try

OpenStudy (anonymous):

somehow, i got 9x=y